OpenStudy (anonymous):

solve the missing variable c = 2rrr , circumference of a circle , c= 7.04 , r = 3.14

OpenStudy (anonymous):

are those three r's? if so you are now missing a variable

OpenStudy (anonymous):

no thats not r,

OpenStudy (anonymous):

is it \[2 \pi r\]?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[C = 2 \pi r\]\[c = 7.04\]\[\pi = 3.14\] plug in everything we know \[7.04=2(3.14)r\]\[\frac{7.04}{2*3.14} = r\]